OpenStudy (anonymous):

The cost of producing pots is the initial investment, $175, plus $3.50 in materials per pot. Therefore, the function C(p) = 175 + 3.5p defines the cost, C, of producing a number of pots, p. To the nearest cent, how much will it cost to produce 125 pots? (Enter only the number without the dollar sign, such as 25.03 for twenty-five dollars and 3 cents.) i got 612.50 then follow up question asked How many pots can be produced if the budget is limited to $450? (Enter only the number.) i got 128 but i got it wrong can anyone help me?

OpenStudy (saifoo.khan):

congrats, u r correct.

OpenStudy (anonymous):

i dont get it, it told me the 128 is wrong

OpenStudy (saifoo.khan):

Lol , i said about 612.

OpenStudy (anonymous):

450 = 175 + 3.5p solve for p.

OpenStudy (saifoo.khan):

Smooth math, please post your problem on left

OpenStudy (anonymous):

Um... I'm not asking a question.

OpenStudy (anonymous):

I'm proposing a method of solving.

OpenStudy (saifoo.khan):

LOl, Sorry!

OpenStudy (anonymous):

Haha it's fine. Kind of funny actually.

OpenStudy (anonymous):

But yeah, you have 450 dollars to spend and the cost is represented in terms of p by C(p) = 175 +3.5p. Set this equal to 450 and solve for p. This gives 450 = 175 + 3.5p 275 = 3.5p p = 275/3.5

OpenStudy (saifoo.khan):

78.57

OpenStudy (anonymous):

Naturally you can't make .57 of a pot, so you only make 78.

OpenStudy (anonymous):

when i solve for p is about 78.57

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

My pleasure.